Eaton Tool Company has fixed costs of $525,000, sells its units for $106, and has variable costs of $56 per unit.

a. Compute the break-even point.
  



b. Ms. Eaton comes up with a new plan to cut fixed costs to $400,000. However, more labor will now be required, which will increase variable costs per unit to $59. The sales price will remain at $106. What is the new break-even point? (Round your answer to the nearest whole number.)
  



c. Under the new plan, what is likely to happen to profitability at very high volume levels (compared to the old plan)?
  

  • Profitability will be less

  • Profitability will be more
Answer #1

a. Let the break even point be X, then  

Total Expense = Total Revenue

525000 + 56*X = 106*X

50*X = 525000

X = 10500 Units

b. Let the break even point be Y, then

400000 + 59*Y = 106*Y

47*Y = 400000

Y = 8511 Units

c. Profitability will be less

Under the new plan, profitability will be less at very high volumes as the variable cost has increased compared to the old.plan.